How many grams of NaOH (molar mass = 40.000 g/mol) is required to prepare 100.0mL of 0.125M solution?
Answer #1

Molarity=0.125M(mol/L)

Volume=0.1L

Mole=0.125*0.1= 0.0125 mol

Mass of naoh= 0.0125*40= 0.5 gram naoh